# LLM-FE: Automated Feature Engineering for Tabular Data with LLMs as Evolutionary Optimizers

### v1

@@ 6.1 Ablation Study
We perform an ablation study on the classification datasets listed in Table 2 to assess the contribution of each component in LLM-FE. Figure 2 illustrates the impact of individual components on overall performance, using XGBoost for prediction and GPT-3.5-Turbo as the LLM backbone. We report the accuracy aggregated over all the classification datasets. In the w/o Domain Knowledge’ setting, dataset and task-specific details are removed from the prompt to assess the impact of domain knowledge on performance. Feature names are anonymized and replaced with generic placeholders such as C1, C2, $\dots$, Cn, effectively removing any semantic meaning that could provide contextual insights about the problem. Without domain knowledge, the classification performance significantly drops to 0.838, underscoring its critical role in generating meaningful features. Removing the evolutionary refinement (w/o Evolutionary Refinement’ setting) also leads to a considerable decline in performance, emphasizing the importance of iterative data-driven feedback in addition to domain knowledge for refining hypotheses. Lastly, ablation results show that eliminating data examples from the LLM prompt (w/o Data Examples’ variant) leads to only a slight performance drop across datasets. This suggests that providing data examples in the prompt offers limited benefit, as LLMs struggle to comprehend the nuances and patterns within the data samples. The best-performing variant (LLM-FE) demonstrates the positive contribution of each component including domain knowledge, evolutionary search, and example-based guidance in enhancing performance beyond the base model.
6.2 Qualitative Analysis
Impact of Evolutionary Refinement.
Impact of Domain Knowledge.
Figure 4 highlights the qualitative and quantitative benefits of domain-specific feature transforms. We demonstrate this using two datasets: the Breast-W dataset, which focuses on distinguishing between benign and malignant tumors, and the Heart dataset, which predicts cardiovascular disease risk based on patient attributes. These tasks underscore the crucial role of domain knowledge in identifying meaningful features. Using embedded domain knowledge, LLM-FE not only significantly improves accuracy but also provides the reasoning for choosing the given feature, leading to more interpretable feature engineering. For example, in the Heart dataset, LLM-FE suggests the feature ‘Log_Cholesterol’, recognizing cholesterol’s critical role in heart health and applying a logarithmic transformation to reduce the impact of outliers and stabilize the variance. In contrast, the ‘w/o Domain Knowledge’ variant arbitrarily combines existing features, leading to uninterpretable transformations and reduced overall performance (Figure 4 (a)). Similarly, for breast cancer prediction, LLM-FE identifies ‘proliferation_activity’ a biologically relevant metric leading to performance improvement, whereas the absence of domain knowledge results in a simple mean of all features, lacking interpretability and clinical significance (Figures 4(b) and 4(c)).
@@ 7 Conclusion
Impact Statement
The introduction of LLM-FE as a framework for leveraging LLMs in automated feature engineering has the potential to significantly impact the field of machine learning for tabular data by improving predictive performance and reducing the manual effort involved in feature engineering. This is particularly beneficial in resource-intensive domains where efficient feature extraction and transformation are crucial for accelerating model development. By combining domain expertise with evolutionary optimization, LLM-FE improves over existing methods, which often struggle to identify optimal feature transformations. While LLM-FE currently targets feature engineering, its potential could be extended to broader data-centric AI tasks. Future extensions could include automated data cleaning (such as imputation and outlier detection), exploratory data analysis, and data augmentation, further strengthening the quality and robustness of datasets in tabular learning. LLM-FE could also be extended to support model tuning and hyperparameter optimization, contributing to a more streamlined and interpretable machine learning pipeline and improving the generalization of the model.
Reproducibility Statement
To ensure the reproducibility of our work, we provide comprehensive implementation details of LLM-FE. Section 3 outlines the full methodology, while Appendix B.2 offers an in-depth description of the framework, including the specific LLM prompts used. The datasets employed in our experiments are also detailed in Appendix 5. Additionally, we plan to release our code and data at this repository to facilitate further research.
@@ Appendix A Dataset Details
Table 5 describes the diverse collection of datasets spanning three major categories: (1) binary classification, (2) multi-class classification, and (3) regression problems used in our evaluation. The datasets were primarily sourced from established platforms, including OpenML (Vanschoren et al., 2014; Feurer et al., 2021), UCI (Asuncion et al., 2007), and Kaggle. We specifically selected datasets with descriptive feature names, excluding those with merely numerical identifiers. Each dataset includes a task description, enhancing contextual understanding for users. Our selection encompasses not only small datasets but also larger ones, featuring extensive data samples and high-dimensional datasets with over 50 features. This diverse and comprehensive selection of datasets represents a broad spectrum of real-world scenarios, varying in both feature dimensionality and sample size, thereby providing a robust framework for evaluating feature engineering works.
@@ B.1 Baselines
FeatLLM uses an LLM to generate rules to binarize features that are then used as input to a simple model such as linear regression. We adapt the open-source featllm [^4] https://github.com/Sungwon-Han/FeatLLM implementation, modifying the pipeline to use an XGBoost model for inference. To ensure a fair comparison with other methods, we provide the entire training dataset to train the XGBoost model while using only a subset of the dataset (10 samples) to the LLM to generate binary features. We report the results through an ensemble over three samples to maintain consistency with LLM-FE.
@@
We utilize the official implementation of CAAFE, [^5] https://github.com/noahho/CAAFE maintaining all parameter settings as specified in the original repository. Following their workflow, we preprocess the data using their pipeline before inputting them into the prediction model after the feature engineering process.
@@ B.2 LLM-FE
Hypothesis Generation.
Figure 5 presents an example prompt for the balance-scale dataset. The prompt begins with general instructions, followed by dataset-specific details, such as task descriptions, feature descriptions, and a subset of data instances serialized and expressed in natural language. To introduce diversity in prompting, we randomly sample between this approach and an alternative set of instructions, encouraging the LLM to explore a wider range of operators from OpenFE (Zhang et al., 2023), as prior LLMs tend to favor simpler operators (Küken et al., 2024) (see Figure 6). By providing this structured context, the model can leverage its domain knowledge to generate semantically and contextually meaningful hypotheses for new feature optimization programs.
@@
Following the ‘islands’ model used by (Cranmer, 2023; Shojaee et al., 2024; Romera-Paredes et al., 2024), we maintain the generated hypotheses along with their evaluation scores in a memory buffer comprising multiple islands ($m=3$) that evolve independently. Each island is initialized with a basic feature transformation program specific to the dataset. Each island is initialized with a simple feature transformation program specific to the dataset (def modify_features_v0()in Figure 5(d)). In each iteration, novel hypotheses and their validation metrics are incorporated into their respective islands only if they exceed the island’s current best score. Within each island, we additionally cluster feature discovery programs based on their signature, characterized by their validation score. Feature transformation programs that produce identical scores are consolidated together, creating distinct clusters. This clustering approach helps preserve diversity by ensuring that programs with varying performance characteristics remain in the population. We leverage this island model to construct prompts for the LLM. After an initial update of the prompt template with dataset-specific information, we integrate in-context demonstrations from the buffer. Following (Shojaee et al., 2024; Romera-Paredes et al., 2024), we randomly select one of the $m$ available islands. Within the chosen island, we sample $k=2$ programs to serve as in-context examples. To sample programs, we first select clusters based on their signatures using the Boltzmann selection strategy (De La Maza & Tidor, 1992) to sample clusters based on their signatures with a preference for clusters with higher scores. Let $s_{i}$ be the score of the i-th cluster, and probability $P_{i}$ for selecting i-th cluster is given as:
@@
Appendix C Additional Analysis
@@ C.1 Transferability of Generated Features
While traditional approaches typically use the same model for both feature generation and inference, we demonstrate that the features generated by one model can be utilized by other models. Following (Nam et al., 2024), we use XGBoost, a computationally efficient decision tree-based model, to generate features to be used by more complex architectures for inference. As demonstrated in Table 6, XGBoost-generated features show an improvement in the performance of MLP and TabPFN over its base versions. This cross-architecture performance improvement suggests that the generated features capture meaningful data characteristics that are valuable across different modeling paradigms.
Table 6: Comparative analysis of LLM-FE using feature transfer. We use XGBoost to perform feature engineering and apply these features to MLP and TabPFN inference models (indicated as LLM-FE XGB). We report the accuracy for classification tasks and normalized root mean square error for regression tasks. We report the mean values and standard deviation across five random splits. bold: indicates the best performance.
@@
Base 0.745 $\pm$ 0.193 0.871 $\pm$ 0.027
LLM-FEXGB GPT-3.5-Turbo 0.763 $\pm$ 0.216 0.848 $\pm$ 0.017
LLM-FE GPT-3.5-Turbo 0.784 $\pm$ 0.175 0.631 $\pm$ 0.043
@@
Base 0.852 $\pm$ 0.133 0.289 $\pm$ 0.016
LLM-FEXGB GPT-3.5-Turbo 0.861 $\pm$ 0.136 0.287 $\pm$ 0.015
LLM-FE GPT-3.5-Turbo 0.863 $\pm$ 0.132 0.286 $\pm$ 0.015
C.2 Robustness to Noise
C.3 Scalability Analysis
In addition to smaller classification datasets (see Section 5), we evaluate the effectiveness of our method on more complex datasets, including (1) high-dimensional datasets with a large number of features and (2) large-scale datasets with a substantial number of instances. We report the performance of the XGBoost model on these classification datasets using accuracy. For LLM-based methods, we utilize GPT-3.5-Turbo as the LLM backbone. Based on the findings from Section 5, we select the best-performing traditional feature engineering approach (OpenFE) and LLM-based baseline (CAAFE) for comparison with LLM-FE.
Table 7: Performance of XGBoost on complex datasets using various Feature Engineering Methods on (1) High-dimensional classification datasets, and (2) Large-scale classification datasets, evaluated using accuracy (higher values indicate better performance). We report the mean values and standard deviation across five random splits. : denotes failure for all runs due to execution errors. bold: indicates the best performance. underline: indicates the second-best performance. ‘n’: indicates the number of samples; ‘p’: indicates the number of features.
n p Base OpenFE CAAFE LLM-FE
Large scale datasets
adult 48.8k 14 0.873 $\pm$ 0.002 0.873 $\pm$ 0.002 0.872 $\pm$ 0.002 0.874 $\pm$ 0.003
bank-marketing 45.2k 16 0.906 $\pm$ 0.003 0.908 $\pm$ 0.002 0.907 $\pm$ 0.002 0.907 $\pm$ 0.002
cdc diabetes 253k 21 0.849 $\pm$ 0.001 0.849 $\pm$ 0.001 0.849 $\pm$ 0.001 0.849 $\pm$ 0.001
jungle_chess 44.8k 6 0.869 $\pm$ 0.001 0.900 $\pm$ 0.004 0.901 $\pm$ 0.038 0.969 $\pm$ 0.004
High dimensional datasets
covtype 581k 54 0.870 $\pm$ 0.001 0.885 $\pm$ 0.007 0.872 $\pm$ 0.003 0.882 $\pm$ 0.003
communities 1.9k 103 0.706 $\pm$ 0.016 0.704 $\pm$ 0.009 0.707 $\pm$ 0.013 0.711 $\pm$ 0.012
arrhythmia 452 279 0.657 $\pm$ 0.019 0.659 $\pm$ 0.018
myocardial 1.7k 111 0.784 $\pm$ 0.023 0.787 $\pm$ 0.026 0.789 $\pm$ 0.023 0.789 $\pm$ 0.023
Mean Rank 2.13 2.00 1.25
Analysis over high-dimensional dataset.
High-dimensional data in machine learning presents unique challenges due to the ‘curse of dimensionality’, where an increase in the feature space leads to sparse data distributions. This sparsity can degrade model performance and complicate feature engineering, making it difficult to identify relevant features and generate new ones to support the downstream task. To evaluate high-dimensional datasets, we focus on datasets containing more than 50 features as part of our analysis. As shown in Table 7, the general trend is that model performance gradually degrades as the number of features increases, though there are exceptions. Our LLM-FE framework addresses these challenges by carefully balancing dimensionality reduction with the retention of meaningful information, ensuring robust model generalization. Notably, our framework outperforms all feature engineering baselines for XGBoost.
Analysis over large-scale dataset.
Large-scale datasets present unique challenges compared to high-dimensional data. While a large number of samples can help reduce overfitting, the sheer volume of data can diminish the ability to identify and extract meaningful insights. Furthermore, even when meaningful features are identifiable, their impact may be reduced due to the dataset’s size. Features that perform well on smaller subsets may struggle to generalize across the entire distribution. Table 7 illustrates these challenges, showing how the presence of a large number of samples can obscure underlying patterns, making improvements from feature engineering negligible. Despite these challenges, LLM-FE shows better results compared to other baseline methods.
C.4 More Details About the Results
@@
We extend the results from Section 5, showcasing the performance improvements achieved by LLM-FE across various prediction models. Specifically, we employ XGBoost, MLP, and TabPFN to generate features and subsequently use the same models for inference. As shown in Table 8, the features using GPT-3.5-Turbo by LLM-FE consistently enhance model performance across different datasets, outperforming the base versions trained without feature engineering.

### v2

@@ 6.1 Ablation Study
Figure 2: Aggregated ablation study results across classification datasets, showcasing the impact of individual components on LLM-FE’s performance: (a) Data Examples, (b) Domain Knowledge, and (c) Evolutionary Refinement. Values are normalized with respect to the base LLM-FE model to facilitate fair comparison across conditions.
@@
We perform an ablation study on the classification datasets (<10,000 samples) listed in Table 2 to assess the contribution of each component in LLM-FE. Figure 2 illustrates the impact of individual components on overall performance, using XGBoost and GPT-3.5-Turbo. We report the accuracy aggregated and normalized over all the datasets. In the w/o Domain Knowledge’ setting, dataset and task-specific details are removed from the prompt and feature names are anonymized with generic placeholders such as C1, C2$\dots$, Cn. In this way, we remove any semantic meaning that could provide contextual insights about the problem. Without domain knowledge, the performance significantly drops to 0.626, underscoring its critical role in generating meaningful features. The w/o Evolutionary Refinement’ setting) also leads to the greatest decline in performance (0.587), emphasizing the importance of iterative data-driven feedback in addition to domain knowledge for refining feature transforms. Lastly, the results show that w/o Data Examples’ variant leads to only a slight performance drop, as LLMs might struggle to comprehend the nuances and patterns within the data samples. LLM-FE benefits significantly from each component, leading to an improved performance.
Figure 3: Qualitative Analysis on Impact of Domain Knowledge. illustrating how LLM-FE (b) utilizes domain knowledge to create meaningful features with descriptions, in contrast to feature engineering without domain insights (a) leading to uninterpretable outputs.
Figure 4: Quantitative impact of domain knowledge on model accuracy. Using domain knowledge boosts performance compared to both the base model and LLM-FE without domain knowledge.
Figure 5: Performance Trajectory Analysis. for LLM-FE w/o evolutionary refinement and LLM-FE. LLM-FE demonstrates a better trajectory, highlighting the advantage of evolutionary refinement.
6.2 Impact of Domain Knowledge and Evolutionary Refinement
Figure 3 illustrates the qualitative benefits of incorporating domain knowledge into feature engineering. In this example, two approaches are contrasted: one without domain knowledge (Figure 3(a)), and LLM-FE guided by domain-specific insights through an LLM-based feature engineering (Figure 3(b)). The domain-agnostic variant creates arbitrary transformations, such as combining features C1 and C3 using a square root of their product and dropping feature C2 without clear justification. In contrast, LLM-FE leverages its embedded knowledge to derive interpretable and clinically meaningful features. Figure 5 presents a quantitative comparison of model performance on the same dataset, showing that LLM-FE with domain knowledge achieves the highest accuracy, outperforming both the base model and LLM-FE without domain knowledge. Figure 5 illustrates the validation accuracy trajectory of LLM-FE with and without evolutionary refinement across 20 iterations. The variant without refinement shows early improvement but quickly plateaus, indicating convergence to a local optimum. In contrast, LLM-FE continues to improve across iterations, achieving higher accuracy overall. This comparison highlights the effectiveness of evolutionary refinement in enhancing performance by enabling the model to escape local optima and optimize more effectively. For detailed analysis on more datasets, feature quality, and computational efficiency, see Appendix D.
@@ Appendix A Dataset Details
Table 5 describes the diverse collection of datasets spanning three major categories: (1) binary classification, (2) multi-class classification, and (3) regression problems used in our evaluation. The datasets were primarily sourced from established platforms, including OpenML [37, 12], UCI [3], and Kaggle. We specifically selected datasets with descriptive feature names, excluding those with merely numerical identifiers. Each dataset includes a task description, enhancing contextual understanding for users. Our selection encompasses not only small datasets but also larger ones, featuring extensive data samples and high-dimensional datasets with over 50 features. This diverse and comprehensive selection of datasets represents a broad spectrum of real-world scenarios, varying in both feature dimensionality and sample size, thereby providing a robust framework for evaluating feature engineering works.
@@ B.1 Baselines
FeatLLM uses an LLM to generate rules to binarize features that are then used as input to a simple model, such as linear regression. We adapt the open-source featllm [^3] https://github.com/Sungwon-Han/FeatLLM implementation, modifying the pipeline to use an XGBoost model for inference. To ensure a fair comparison with other methods, we provide the entire training dataset to train the XGBoost model while using only a subset of the dataset (10 samples) to the LLM to generate binary features. We report the results through an ensemble over three samples to maintain consistency with LLM-FE.
@@
We utilize the official implementation of CAAFE, [^4] https://github.com/noahho/CAAFE, maintaining all parameter settings as specified in the original repository. Following their workflow, we preprocess the data using their pipeline before inputting it into the prediction model after the feature engineering process.
OCTree.
The official OCTree implementation [^5] https://github.com/jaehyun513/OCTree was modified to keep the data loading and model initialization part common. We implemented OCTree only for classification datasets, as the official implementation is limited to classification datasets, and running for regression datasets on our own could have resulted in incorrect implementation.
@@ B.2 LLM-FE
Feature Generation.
Figure 6 presents an example prompt for the balance-scale dataset. The prompt begins with general instructions, followed by dataset-specific details, such as task descriptions, feature descriptions, and a subset of data instances serialized and expressed in natural language. To introduce diversity in prompting, we randomly sample between this approach and an alternative set of instructions, encouraging the LLM to explore a wider range of operators from OpenFE [45], as prior LLMs tend to favor simpler operators [25] (see Figure 7). The quality of features generated has been detailed in Appendix D.2. By providing this structured context, the model can leverage its domain knowledge to generate semantically and contextually meaningful hypotheses for new feature optimization programs.
@@
Following the ‘islands’ model used by [7, 36, 35], we maintain the generated hypotheses along with their evaluation scores in a memory buffer comprising multiple islands ($m=3$) that evolve independently. Each island is initialized with a basic feature transformation program specific to the dataset. Each island is initialized with a simple feature transformation program specific to the dataset (def modify_features_v0()in Figure 6(d)). In each iteration, novel hypotheses and their validation metrics are incorporated into their respective islands only if they exceed the island’s current best score. Within each island, we additionally cluster feature discovery programs based on their signature, characterized by their validation score. Feature transformation programs that produce identical scores are consolidated together, creating distinct clusters. This clustering approach helps preserve diversity by ensuring that programs with varying performance characteristics remain in the population. We leverage this island model to construct prompts for the LLM. After an initial update of the prompt template with dataset-specific information, we integrate in-context demonstrations from the buffer. Following [36, 35], we randomly select one of the $m$ available islands. Within the chosen island, we sample $k=2$ programs to serve as in-context examples. To sample programs, we first select clusters based on their signatures using the Boltzmann selection strategy [8] to sample clusters based on their signatures with a preference for clusters with higher scores. Let $s_{i}$ be the score of the i-th cluster, and the probability $P_{i}$ for selecting the i-th cluster is given as:
@@
Appendix C Additional Results
@@ C.1 Transferability of Generated Features
While traditional approaches typically use the same model for both feature generation and inference, we demonstrate that the features generated by one model can be utilized by other models. Following [31], we use XGBoost, a computationally efficient decision tree-based model, to generate features to be used by more complex architectures for inference. As demonstrated in Table 6, XGBoost-generated features show an improvement in the performance of MLP and TabPFN over their base versions. This cross-architecture performance improvement suggests that the generated features capture meaningful data characteristics that are valuable across different modeling paradigms.
Table 6: Comparative analysis of LLM-FE using feature transfer. We use XGBoost to perform feature engineering and apply these features to MLP and TabPFN (indicated as LLM-FE XGB). We report the accuracy for classification tasks and the normalized root mean square error for regression tasks. We report the mean and standard deviation across five random splits. bold: indicates the best performance.
@@
Base 0.745 $\pm$ 0.034 0.871 $\pm$ 0.027
LLM-FEXGB GPT-3.5-Turbo 0.763 $\pm$ 0.030 0.848 $\pm$ 0.017
LLM-FE GPT-3.5-Turbo 0.791 $\pm$ 0.029 0.631 $\pm$ 0.043
@@
Base 0.852 $\pm$ 0.028 0.289 $\pm$ 0.016
LLM-FEXGB GPT-3.5-Turbo 0.861 $\pm$ 0.017 0.287 $\pm$ 0.015
LLM-FE GPT-3.5-Turbo 0.863 $\pm$ 0.018 0.286 $\pm$ 0.015
C.2 LLM-FE and Hyperparameter Optimization (HPO)
To assess hyperparameter optimization (HPO) effectiveness, we establish a comprehensive search space for our baseline models and employ the Optuna library [1] for tuning. Our methodology incorporates 400 trials with random sampling across multiple dataset splits. We concentrate our analysis on five classification datasets where baseline models achieve accuracies below 0.8. Table 7 demonstrates that HPO consistently enhances performance across all evaluated datasets for the Base model. Significantly, our proposed method LLM-FE delivers additional performance improvements even in the HPO-optimized setting. These findings confirm that while HPO provides meaningful performance gains, LLM-FE offers complementary and substantial enhancements, reinforcing its effectiveness as a feature engineering approach independent of hyperparameter tuning benefits.
Table 7: Comparison of XGBoost classification accuracy across datasets using Base and LLM-FE models, evaluated under (a) without hyperparameter optimization (HPO) and (b) with HPO.
Dataset w/o HPO w/ HPO
Base LLM-FE Base LLM-FE
eucalyptus 0.655 $\pm$ 0.024 0.668 $\pm$ 0.027 0.659 $\pm$ 0.022 0.678 $\pm$ 0.020
credit-g 0.751 $\pm$ 0.019 0.766 $\pm$ 0.025 0.761 $\pm$ 0.022 0.784 $\pm$ 0.017
cmc 0.528 $\pm$ 0.030 0.531 $\pm$ 0.015 0.554 $\pm$ 0.026 0.578 $\pm$ 0.021
blood-transfusion 0.674 $\pm$ 0.017 0.782 $\pm$ 0.017 0.777 $\pm$ 0.021 0.805 $\pm$ 0.009
vehicle 0.754 $\pm$ 0.016 0.761 $\pm$ 0.027 0.776 $\pm$ 0.035 0.801 $\pm$ 0.033
C.3 Additional Results
@@
We extend the results from Section 5, showcasing the performance improvements achieved by LLM-FE across various prediction models. Specifically, we employ XGBoost, MLP, and TabPFN to generate features and subsequently use the same models for inference. As shown in Table 8, the features using GPT-3.5-Turbo by LLM-FE consistently enhance model performance across different datasets, outperforming the base versions trained without feature engineering. To further assess the generalizability of LLM-FE, we conducted experiments on smaller prediction models like CatBoost and Logistic Regression. From Table 9 that LLM-FE outperforms the respective base models for most of the datasets.
@@
Table 9: Performance improvement with LLM-FE on CatBoost and Logistic Regression. We report the mean and standard deviation over five splits. We use Accuracy for classification datasets, with a higher value indicating better performance. bold: indicates the best performance.
Dataset Logistic Regression CatBoost
Base LLM-FE Base LLM-FE
balance-scale 0.874 $\pm$ 0.024 0.994 $\pm$ 0.012 0.870 $\pm$ 0.020 1.000 $\pm$ 0.000
breast-w 0.955 $\pm$ 0.014 0.962 $\pm$ 0.008 0.957 $\pm$ 0.009 0.962 $\pm$ 0.008
blood-transfusion 0.799 $\pm$ 0.014 0.799 $\pm$ 0.009 0.742 $\pm$ 0.012 0.751 $\pm$ 0.036
car 0.690 $\pm$ 0.017 0.696 $\pm$ 0.031 0.999 $\pm$ 0.001 0.999 $\pm$ 0.001
cmc 0.520 $\pm$ 0.019 0.525 $\pm$ 0.012 0.518 $\pm$ 0.028 0.548 $\pm$ 0.027
credit-g 0.764 $\pm$ 0.006 0.780 $\pm$ 0.015 0.714 $\pm$ 0.046 0.700 $\pm$ 0.021
eucalyptus 0.671 $\pm$ 0.036 0.667 $\pm$ 0.042 0.436 $\pm$ 0.027 0.509 $\pm$ 0.050
heart 0.877 $\pm$ 0.021 0.872 $\pm$ 0.025 0.845 $\pm$ 0.015 0.839 $\pm$ 0.018
pc1 0.931 $\pm$ 0.003 0.935 $\pm$ 0.003 0.929 $\pm$ 0.005 0.932 $\pm$ 0.012
tic-tac-toe 0.997 $\pm$ 0.004 0.996 $\pm$ 0.005 0.673 $\pm$ 0.027 0.747 $\pm$ 0.152
vehicle 0.772 $\pm$ 0.028 0.769 $\pm$ 0.015 0.719 $\pm$ 0.045 0.725 $\pm$ 0.033
C.4 Robustness to Noise
Appendix D Qualitative Analysis
D.1 Computational Efficiency
Figure 9: Pareto Plot: comparing trade-off between performance (accuracy) vs time (in seconds) for LLM-FE and other feature engineering baselines.
Automated feature engineering methods, both classical and LLM-based, universally employ model training and validation to evaluate feature relevance. This evaluation strategy represents standard methodology across all automated feature engineering approaches rather than an additional computational burden specific to LLM-FE. We conduct our efficiency-performance trade-off analysis on the datasets with higher sample counts from Section 5, as these datasets present greater complexity with their substantial number of samples and features. Our comparative Pareto analysis (Figure 9) presents the base model alongside various feature engineering baselines. Our proposed method, LLM-FE, demonstrates Pareto optimality by achieving superior performance with substantially reduced computational requirements compared to existing methods, which either exhibit longer execution times or yield inferior performance metrics on these datasets. Only the base method requires less computation time than LLM-FE, but at a significant performance cost. This positions LLM-FE as the optimal solution in the efficiency-performance space, delivering state-of-the-art results with reasonable computational demands even when handling datasets of considerable complexity.
D.2 Complexity of Generated Features
Despite a tendency of LLMs to prefer simpler operators [25] (e.g., addition, subtraction, absolute value), resulting in a less diverse feature set compared to specialized feature-engineering tools, LLM-FE demonstrates the valuable capability of identifying certain complex and informative transformations that are rarely generated by conventional LLM-based automated methods. Specifically, complex operators like groupbythenmean, groupbythenmin, groupbythenmax, residual, and sigmoid are also recommended frequently by LLM-FE, as illustrated in Figure 10. Such complex operations have the potential to capture meaningful patterns involving group-based aggregation that simpler transformations may miss. Thus, while further refinement is needed to balance operator selection, the ability of LLM-FE to discover nuanced, aggregation-based features emphasizes its promising role as a complementary technique in the broader automated feature engineering toolkit.
Figure 10: Frequency of Feature Engineering Operators. We compare the operators for LLM-FE with simple LLM-based methods.
D.3 Impact of Domain Knowledge
Figure 11 highlights the qualitative and quantitative benefits of domain-specific feature transforms. We demonstrate this using two datasets: the Breast-W dataset, which focuses on distinguishing between benign and malignant tumors, and the Heart dataset, which predicts cardiovascular disease risk based on patient attributes. These tasks underscore the crucial role of domain knowledge in identifying meaningful features. Using embedded domain knowledge, LLM-FE not only significantly improves accuracy but also provides the reasoning for choosing the given feature, leading to more interpretable feature engineering. For example, in the Heart dataset, LLM-FE suggests the feature ‘Log_Cholesterol’, recognizing cholesterol’s critical role in heart health and applying a logarithmic transformation to reduce the impact of outliers and stabilize the variance. In contrast, the ‘w/o Domain Knowledge’ variant arbitrarily combines existing features, leading to uninterpretable transformations and reduced overall performance (Figure 11(a)). Similarly, for breast cancer prediction, LLM-FE identifies ‘proliferation_activity’ a biologically relevant metric leading to performance improvement, whereas the absence of domain knowledge results in a simple mean of all features, lacking interpretability and clinical significance (Figures 11(b) and 11(c)).
D.4 Impact of Evolutionary Refinement